Sunrise​ Manufacturing, Inc. Sunrise​ Manufacturing, Inc., a U.S. multinational​ company, has the following debt components in its consolidated capital​ section,

Tax rate 40%

25-year US$ bonds 30,000,000

5-year US$ bonds 6,000,000

10-year eurobonds (euros) 8,000,000

20-year yen bonds (yen) 950,000,000

Before-tax cost of 25-year US$ bonds 7.0%

Before-tax cost of 5-year US$ bonds 3.5%

Before-tax cost of 10-year eurobonds 4.5%

Before-tax cost of 20-year yen bonds 2.5%

Spot rate ($/euro) 1.2900

Spot rate ($/pound) 1.8900

Spot rate (yen/$) 107.00 ​

Sunrise's shareholders' equity is ​$55,000,000 and its finance staff estimates their cost of equity to be 17.5​%. Current exchange rates are also listed in the table. Income taxes are 40​% around the world after allowing for credits. Calculate​ Sunrise's weighted average cost of capital. Are any assumptions implicit in your​ calculation?

What is​ Sunrise's weighted average cost of​ capital? (Round two decimal places.)
Answer #1

The 10 year Euro bonds in USD would be = 8,000,000 *1.29 = 10,320,000 USD

The 20 year Yen Bonds in USD would be = 950,000,000/107 = 8,878,504.67 USD

Total debt =30,000,000 + 6,000,000 +10,320,000 + 8,878,504.67 = 55,198,505 USD

Cost of debt before tax = 30/55.198505 * 7% +6/55.198505 * 3.5% + 10.32/55.198505 *4.5% + 8.878505/55.198505 * 2.5 % = 5.4283%

After tax cost of debt = 5.4283*(1-0.4) = 3.2570%

Debt = 55,198,505 USD, Equity = 55,000,000 USD

Total =110,198,505 USD

WACC = 55,198,505/110,198,505 * 3.2570% + 55,000,000/110,198,505 *17.5% = 0.1037 = 10.37%

Sunrise's weighted average cost of​ capital = 10.37%